OpenStudy (abbycross167):

Can someone please show me the steps to answer this question? Find x when 0.02x + 0.7 > 0.8

OpenStudy (kamibug):

Subtract 0.7 from both sides then divide both sides by 0.02. :)

OpenStudy (abbycross167):

could you show me how to do that?

OpenStudy (abbycross167):

0.02-0.7x + 0.7 > 0.8-0.7 IS this what you meant?

OpenStudy (kamibug):

Close...the x has to stay on 0.02 0.02x + 0.7 - 0.7 > 0.8 - 0.7

OpenStudy (abbycross167):

oh ok

OpenStudy (mathmale):

Please share what you currently have.

OpenStudy (abbycross167):

so then it'll be 0.02x+0.7-0.7>0.8-0.7 = 0.02x+>0.1?

OpenStudy (mathmale):

Yes, good! Now, as a final step, you must solve for x (isolate x). How would you go about doing that?

OpenStudy (abbycross167):

wait wait @KamiBug when do i need to divide ?

OpenStudy (kamibug):

To isolate x as @mathmale said, you'd divide now that you have 0.02x > 0.1 \[\frac{ 0.02x }{ 0.02 } > \frac{ 0.1 }{ 0.02 }\]

OpenStudy (abbycross167):

So 0.02x/0.02>0.1/0.02= 1>5?

OpenStudy (mathmale):

0.02x+>0.1 should actually be:\[0.02x \ge 0.1\]

OpenStudy (mathmale):

Divide each side of this inequality by 0.02. That will give you x alone on the left side. Please share your work.

OpenStudy (abbycross167):

wait i'm a little confused

OpenStudy (abbycross167):

wouldn't it be 1>5?

OpenStudy (mathmale):

No. Sorry the stuff I typed earlier didn't come out clearly.\[0.02 \ge0.1\]

OpenStudy (mathmale):

is to be divided on both sides by 0.02. Show your work, please.

OpenStudy (abbycross167):

so 0.02/0.02=>0.1/0.2?

OpenStudy (abbycross167):

@KamiBug

OpenStudy (kamibug):

Sorry I'm late :o So right now we have 0.02x > 0.1 Divide bith sides by 0.02 like this 0.02x / 0.02 > 0.1 / 0.02 On the left side, the 0.02 cancles out and you're left only with x. x > 0.1 / 0.02 Finish simplifying... :)

OpenStudy (abbycross167):

>5?

OpenStudy (abbycross167):

thank you very much!! : )